Fermi National Laboratory
 

Accelerator Update for January 12, 2009 - January 14, 2009

 

Monday Jan 12

The day shift began with Operations monitoring the Tevatron (TeV) store 6718 (162.54E30) and the Recycler stash (140.9E10), with the Antiproton Source stacking (33.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE and NuMI experiments taking beam.

Operations reported at 10:01 AM that experts had begun work on replacing the Booster LCW pond water pump and on a MI RF station’s (MIRF10) bias power supply. These jobs caused about 20 minutes of downtime.

At 11:42 AM, a Linac Klystron RF station (KRF5) tripped off for the first of four times in about 2.5 hours. It reset.

Time Line Generator (TLG) experts began a study at 2:20 PM in preparation for the multibatch proton injections for a shot setup on Tuesday, 1/13/09.

At 3:04 PM, KRF7 tripped off. It reset.

At 3:39 PM, store 6718 was lost when the TeV suffered a sector D4 quench. TeV experts investigated.

TeV experts reported at 3:51 PM that the quench was due to the TLG studies.

Operations reported at 4:06 PM that the Recycler had stashed antiprotons seven times during the day shift.

At 4:32 PM, LRF5 tripped off. It reset.

Operations put the TeV through a dry squeeze at 6:14 PM.

Operations put the TeV through the first of two wet squeezes beginning at 6:54 PM.

Operations put the TeV into shot setup at 7:52 PM.

At 8:38 PM, LRF2 tripped off. It reset.

At 8:43 PM, the Tevatron Electron Lens (TEL1) tripped off on a vacuum interlock indication. It reset.

Operations established store 6722 at 9:43 PM with an initial luminosity of 347.8E30.

At 11:15 PM, KRF5 tripped off. It reset.

Operations reported at midnight that the Recycler had stashed six times during the evening shift.

Tuesday Jan 13

The midnight shift began with Operations monitoring store 6722 (211.77E30) and the stash (105.9E10), with Pbar stacking (19.1E10), and with NuMI and MiniBooNE taking beam.

At 12:29 AM, BRF7 tripped off. It reset.

At 12:57 AM, KRF5 tripped off. It reset.

At 4:41 AM, BRF8 tripped off. It reset.

At 6:45 AM, LRF2 tripped off. It reset.

At 6:56 AM, LRF4 tripped off. It reset.

Operations reported at 8 AM that the Recycler had stashed 10 times during the midnight shift.

At 8:15 AM, a Linac quadrupole power supply (L:QPS111) tripped its breaker. It reset.

Operations informed LCW experts at 8:19 AM that a MI-20 LCW pump had tripped off. They said they would take care of bringing it back online.

At 8:57 AM, KRF5 tripped off. It reset.

At 9:44 AM, Operations reported that KRF5 was continuing to trip off. A controls expert installed some diagnostics.

Around 11:08 AM, the I- Source sparked and caused various Booster RF stations to trip off. Everything reset.

At 11:21 AM, a Linac quadrupole power supply (QPS111) tripped off. It reset.

At 12:41 PM, KRF5 tripped off. It reset.

Operations terminated store 6722 at 12:54 PM.

Operations put the TeV into shot setup at 1:06 PM.

At 1:39 PM, the TeV aborted while operators were investigating a cryo system tank farm pressure alarm.

At 2 PM, the TeV aborted again.

At 2:01 PM, KRF7 tripped off. It reset.

Operations reported at 2:18 PM that due to the tank farm alarm and to an unstable wet engine at TeV sector A4, operators were not going to load antiprotons into the TeV. Cryo system experts investigated the A4 wet engine problems.

Cryo experts reported at 2:37 PM that all was well with the A4 wet engine. Operations resumed shot setup.

Operations turned off the Linac at 2:42 PM to make an access. Preaccelerator experts will also access the I- Source dome.


At 3:07 PM, the TeV suffered a quench at sectors A1 and C2. TeV experts investigated.

The Preaccelerator and Linac came out of access at 3:19 PM.

Electron Cooling (ECOOL) experts accessed the Pelletron at 3:26 PM.

Operations reported at 4:05 PM that the Recycler had stashed five times during the day shift.

At 4:19 PM, Linac vacuum valves (VLNLAM and VLNDMP) closed. An expert investigated. Linac resumed sending beam about 10 minutes later.

At 4:34 PM, Operations reported that ECOOL had tripped off about 40 minutes earlier. It reset.

At 5:17 PM, Operators reported that they were replacing a failed power supply (T:QF28). They did not believe this was the cause of the TeV quench.

At 6:08 PM, Operations reported that the TeV sector A1 cryo system was warming up. The Cryo Coordinator reported a two-phase-to-cryostat vacuum leak. He thinks that the leak can be overcome with helium without a TeV access.

Operations reported at 7:45 PM that the A1 cryo system was slowly coming around.

The Cryo Coordinator reported at 8:01 PM that the A1 cryo system was good to go.

At 8:25 PM, KRF5 tripped off. It reset.

Operations put the TeV through a dry squeeze at 8:39 PM. TeV experts began a study period.

At 9:17 PM, Operations found a problem with a Pbar power supply (Q107). Operations called in an expert.

At 9:50 PM LRF4 tripped off. It reset.

At 10 PM, BRF18 began tripping off. Operators bypassed the station and brought BRF19 online in its place.

At 10:23 PM, Operations reported that Pbar Q107 work was continuing. There was no stacking due to this problem.

Operations reported at midnight that the Recycler had stashed six times during the evening shift.

Wednesday Jan 14

The midnight shift began with Operations monitoring the stash (179.0E10) and the stack (20.9E10), and with NuMI and MiniBooNE taking beam.

TeV experts completed their study at 12:37 AM.

Operations put the TeV into shot setup at 12:43 AM.

Pbar power supply experts reported at 1:20 AM that they had replaced a 15-volt power supply that was causing poor regulation in Q107.

Pbar resumed stacking at 1:38 AM.

Operations established store 6727 at 2:08 AM with an initial luminosity of 180.97E30.

At 3:43 AM, two Linac vacuum pumps (IPRF17 and IPBD2) tripped off. IPRF17 would reset, but IPBD2 would not. Operations called in experts.

Operations put Pbar into standby mode at 4:03 AM.

Operations prepared Linac for access at 5:31 AM.

At 7:04 AM, the NuMI horn tripped off on a LCW indication. Experts investigated. Operations prepared MI-65 for an access.

A Booster expert reported at 7:50 AM that the BRF18 problem was in the tunnel.

Operations reported at 8:02 AM that the Recycler had stashed two times during the midnight shift.

The Plans for Wednesday

The plans for today are to fix the Linac dump vacuum problem.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/14/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory